# Spontaneous emission of a three-level artificial atom in a one-dimensional open waveguide

### v1

@@ Abstract
We study the dynamical and spectral characteristics of a quantum three-level ladder system, interacting with a continuous electromagnetic field in one-dimensional open waveguide. Common realization of such systems is a waveguide QED setup  — a superconducting artificial atom (transmon), coupled to an open microwave transmission line. We derive an analytical solution for spontaneous emission of initially excited atom, and use it to study the probability of state detection and spectral density of output photon states. We find that for strong coupling of transmon to a waveguide emitted photons show correlation in frequency and can have the same energies, even if the three-level system is anharmonic.
waveguide QED, photon-photon correlation, artificial atom
@@ I Introduction
Waveguide quantum electrodynamics is a modern and rapidly developing field of research that studies the interaction of atoms with photons in confined geometries of waveguides Roy17 ; Gu17 ; Sheremet23 ; Astafiev10 . It has a vast spectrum of applications in quantum optics and photonics, as well as in quantum technologies such as quantum communication and quantum information processing. In contrast with cavity QED Raimond01 ; Haroche06 ; Kibis11 ; Savenko12 ; Kibis13 and circuit QED Blais04 ; Wallraff04 ; Blais21 that are focused more on resonant interaction of atoms and nanostructures with single-mode fields in cavities or resonators, waveguide QED describes interaction with a continuum of photon modes in open waveguides.
@@
Three-level systems play an important role in quantum optics and photonics, and have recently received broad attention in waveguide QED as well. Interaction of three-level emitters (3LE) with photon fields can induce some interesting and important effects, such as electromagnetically induced transparency (EIT) Abdumalikov10 , Autler-Townes splitting Sillanpaa09 ; Anisimov11 , two-photon resonance fluorescence Gasparinetti19 ; Ngaha25 ; Gunin25 , giant cross-Kerr effect Hoi13a , photon-state purification Lopes24 , and many others. Since an excited three-level system can emit two photons via spontaneous decay or scattering interaction, it can be potentially used to induce photon-photon correlations. Creation of correlated photon pairs due to spontaneous decay or scattering on a single atom in 1D waveguides was studied in Shen07a ; Shen07b ; Roy11 ; Zheng12 ; Chen12 ; Li15 ; Sathyamoorthy16 ; Gasparinetti17 ; Gao24 .
@@
A typical realization of a 3LE interacting with a continuum of photons in 1D waveguides is superconducting artificial atoms Kjaergaard20 ; Krantz19 , directly or side-coupled to a microwave transmission line Astafiev10 ; Blais21 . The most common type of a three-level system in this playground is a transmon qubit Koch07 . A transmon has a ladder-type structure of levels, meaning that transitions from a lower to a higher level can only happen via a middle level. Although transmons are usually used for quantum information applications as regular qubits, i.e., two-level systems, the third level can have an important practical role in waveguide QED setups Hoi11 ; Lu21 ; Sultanov25 , and, of course, has wide quantum optics applications (see Hoi13b and references from the previous paragraph).
@@
In this work we study the spontaneous emission of an initially excited three-level system, coupled to a 1D open waveguide. In contrast with more traditional tasks like interaction of radiation with single qubits Shen05 ; Chen11 ; Wang11 or qubit chains Fang14 ; Greenberg15 ; Liao15 ; Greenberg21 ; Greenberg23 , that are usually performed in the single-photon limit, here we employ the two-excitation subspace. Our approach is based on the real-space description of photon fields. We derive an analytical solution of the non-stationary Schrödinger equation and find the probability amplitudes for the complete wavefunction of the system. Using this solution, we analyze probabilities to detect certain states of the system, and the spectral density of output two-photon fields. For a weak radiative decay rate we find that each of the two photons has a distinct frequency, corresponding to the two transitions of the 3LE, which is in good agreement with the experimental work Gasparinetti17 . However, if the radiative decay is approximately ten times higher (which is achievable in current setups), we find that frequency correlation appears, which can result in the creation of two identical photons with equal energies. This correlation is more pronounced for low anharmonicity of the 3LE (i.e., when the levels are almost equidistant). We show that depending on the ratio of decay rates from the upper and middle levels, the spectral density for identical photons can have two maxima: on resonance with the lower two levels, or resonant with a smaller frequency which differs from the first frequency by half the anharmonicity parameter.
@@
The paper is organized as follows. In Section II we give a general description of the system under study and derive the main equations for the wavefunction’s amplitudes. In Section III we derive a general analytical solution in real-space. In Section IV we examine the case of a fully-excited three-level atom, both in real-space and in frequency representations. The main results are placed in this section, including analysis of state-detection probabilities, two-photon emission spectra and creation of correlated photon pairs that share the same frequencies. The conclusion of our work, as well as future prospects, are reflected in Section V. Some mathematical calculations are included in the Appendix.
@@ II Real-space description of the transmon-waveguide system
We consider a system of a ladder-type three-level emitter (3LE), located at the point $x=0$ in a one-dimensional waveguide (see Fig. 1a). In the real-space representation Shen07b ; Shen09 and under rotating wave approximation it can be described by the following Hamiltonian (here and later in the text we assume that $\hbar=1$):
@@
Here states $\left|g\right\rangle,\left|e\right\rangle$ and $\left|f\right\rangle$ are ground, first-excited and second-excited states of the transmon. Corresponding resonant frequencies of first and second excited states are $\Omega_{1},\Omega_{2}$, and $V_{1},V_{2}$ are their coupling constants. Energy levels and corresponding frequencies are shown in Fig. 1b. $\upsilon_{g}$ is a group velocity of photons, $\delta(x)$ is a Dirac delta-function, and $a^{\dagger}(x)$ ($a(x)$) is a creation (annihilation) operator in x-representation, which creates (destroys) photon in $x$ position of the waveguide.
Throughout of this paper we will refer to 3LE as transmon or (artificial) atom, but keep in mind that our approach is universal and can be related to other physical realizations of three-level ladder-type systems.
Figure 1: (a) System under study - three-level emitter (3LE), coupled to the center of one-dimensional waveguide at point $x=0$. Coupling of first and second excited states are $V_{1}$ and $V_{2}$, respectfully. Excited atom can emit photons through spontaneous decay, which are propagate only forward due to usage of chiral approximation (see text for details). (b) Energy levels diagram of the bare transmon and corresponding resonant frequencies.
Hamiltonian (1) is written in the so-called continuous limit Blow90 , meaning that instead of discrete summation over $x$ (or over modes $k$), we have continuous integration. In this limit, commutation relation between photon operators is as follows:
@@
hence the photon operators, usually dimensionless for discrete modes, in continuous limit have dimension of inverse square root of coordinate, $a(x)\sim 1/\sqrt{x}$. The dimension of coupling constants $V_{1},V_{2}$, correspondingly, is $V\sim\omega\sqrt{x}$.
The first two terms in (1) correspond to a free-atom and free-field, and the following two terms describe the interaction of photon fields with first and second excited states of transmon, respectively. Note that there is no direct transition between ground and second-excited state $\left|f\right\rangle\leftrightarrow\left|g\right\rangle$. Hamiltonian (1) is practically Janes-Cumming Hamiltonian, extended for three-level system. It conserves number of excitations in the system. In this work we adhere two excitations due to two possible transitions of the transmon.
Since our atom have three levels, wave function should be written in a two-excitation subspace as follows:
@@
where first probability amplitude $\alpha(t)$ corresponds to fully excited transmon in state $\left|f\right\rangle$ and no photons in a waveguide, second amplitude $\beta(x,t)$ corresponds to transmon excited to $\left|e\right\rangle$ state and one photon at position $x$, and final amplitude $\gamma(x_{1},x_{2},t)$ contains non-excited atom and two photons in $x_{1}$ and $x_{2}$ positions. State $\left|0\right\rangle$ designates vacuum Fock state. Coefficient $1/\sqrt{2}$ in last term is a normalization constant for two-photon state in Fock basis.
Because of continuous limit Blow90 , probability amplitudes with photons have dimension of inverse square root of coordinate: $\beta(x,t)\sim 1/\sqrt{x}$, and $\gamma(x_{1},x_{2},t)\sim 1/x$ due to presence of two photons.
We should also note that wavefunction (3), as well as Hamiltonian (1), describe photon propagation only in forward direction. This is so called chiral approximation, and it can be performed by separating Hamiltonian for odd and even modes Shen07b . Chiral, or non-reciprocal, waveguide can be realized experimentally by creating suppression of photon propagation in backward direction Lodahl17 ; Guimond20 . We are mainly interested in the case of spontaneous decay of initially excited 3LE, which should be symmetrical to 3LE placement. So, in order to simplify analytical derivation, we choose to consider only forward propagation.
Using standard non-stationary Schrödinger equation $i\partial_{t}\left|\Psi\right\rangle=H\left|\Psi\right\rangle$ with Hamiltonian (1) and wave function (3), we derive the equations for all probability amplitudes:
@@
In the last two equations we use the fact that photons are fundamentally indistinguishable, which in our case can be stated as $\gamma(x_{1},x_{2},t)=\gamma(x_{2},x_{1},t)$. All further calculations are based on equations (4-6).
@@ III Probability amplitudes in real-space
To find explicit expressions for all three amplitudes of wave function, we use method, incorporated in Mukhopadhyay24 for calculation of quantum Rabi oscillation dynamics and also employed in Lopes24 for photon state-purification via 3LE interaction. We have already used this approach in Chuikin24 for description of stimulated emission in waveguide QED, which is very similar task to that we solve here.
First, let’s define a Fourier transform for amplitudes with photonic part, which allow us to switch from real-space to k-space of frequency modes:
@@
Inverse transformation is done by changing sign of exponents.
Here and further we assume that dispersion is linear, $k=\omega/\upsilon_{g}$, which is well justified if the atom and photon resonant frequencies are far from cut-off frequency of the waveguide Shen07b ; Shen09 .
By using (7) in (6) together with integral representation of delta function as $\delta(x)=1/2\pi\int dk\;e^{ikx}$, we obtain equation for two-photon amplitude in k-space:
$\begin{gathered}\frac{d}{dt}\gamma(k_{1},k_{2},t)+i\upsilon_{g}(k_{1}+k_{2})\gamma(k_{1},k_{2},t)\\ =-\frac{i}{\sqrt{2}\sqrt{2\pi}}V_{1}{e^{-i\Omega_{1}t}}\left[{\beta(k_{2},t)+\beta(k_{1},t)}\right].\end{gathered}$ (8)
@@
where we assume that at initial time $t=0$ no two-photon states are presented. Transforming this expression back to real-space using inverse of (7), we get:
@@
where we introduced a Heaviside step-function, $\theta(x>0)=1,\quad\theta(x<0)=0$. In order to avoid a jump at $x=0$, we apply standard regulation $\theta(x)=1/2$.
Next thing we need is to find $\beta(x,t)$, which then can be used to determine two-photon amplitude (10). Taking $\gamma(x_{1},x_{2},t)$ for $x_{2}=0$ and substitute it in (5), we obtain:
@@
where we define decay rate of first-excited level as $\Gamma_{1}=V_{1}^{2}/\upsilon_{g}$, as well as some additional parameter:
@@
To find explicit expression for $\beta(x,t)$, we use same procedure as for $\gamma(x_{1},x_{2},t)$. Multiplying (11) by $e^{-ikx}/\sqrt{2\pi}$ and integrating over $x$, we can use Fourier transform (7) to rewrite equation (11) in k-space. Then formal solution can be found as:
@@
where $\beta_{0}(k)=\beta(k,0)$ is an amplitude for incident photon with frequency $\omega=k\upsilon_{g}$ at initial time $t=0$ and $F(k,t)$ is a Fourier transform of coefficient (12). Translating (13) back to real-space, we get:
@@
where we introduce frequency between levels $\left|e\right\rangle$ and $\left|f\right\rangle$ as $\Delta_{\Omega}=\Omega_{2}-\Omega_{1}$ (see Fig. 1b). Change of upper limit in last integral is due to step function $\theta\left(x-\upsilon_{g}t^{\prime}\right)$.
Expression (14), aside from amplitude $\alpha(t-x/\upsilon_{g})$, is also contain shifted value $\beta(\upsilon_{g}t^{\prime}-x,t-x/\upsilon_{g})$, so it can’t be considered explicit. However, by performing a corresponding shift in (14), one can show that recursive part will cut off due to Heaviside functions, and remaining part will contain only terms proportional to incident single-photon field $\beta_{0}$ and amplitude of second-excited state $\alpha(t)$. Detailed calculations can be found in Appendix A. Thus, we obtain the explicit expression for a single-photon amplitude:
@@
By taking (15) at point $x=0$, we can easily find the remaining amplitude of a second-excited state from (4):
$\begin{gathered}\alpha(t)=\alpha(0)e^{-\frac{\Gamma_{2}}{2}t}\\ -iV_{2}e^{-\frac{\Gamma_{2}}{2}t}\int\limits_{0}^{t}d\tau e^{\left({i\Delta_{\Omega}-\frac{\Gamma_{1}}{2}+\frac{\Gamma_{2}}{2}}\right)\tau}\beta_{0}(-\upsilon_{g}\tau),\end{gathered}$ (16)
where $\alpha(0)$ corresponds to initial state with fully excited transmon, and $\Gamma_{2}=V_{2}^{2}/\upsilon_{g}$ is a decay rate of second-excited state, similar to $\Gamma_{1}$.
Obtained expressions (10), (15) and (16) are the key results of this section. Taking particular initial state, we can derive explicit analytic expression for full wave function (3). These equations contain two possible initial conditions: 1) Transmon fully excited in state $\left|f\right\rangle$ and there are no free-photons in a waveguide; 2) Transmon is excited to state $\left|e\right\rangle$ and there is a single incident photon with arbitrary line shape $\beta_{0}\equiv\beta(k,0)$. First condition describes spontaneous emission of fully excited three-level system, and we analyze it below in details. Second one describes process of stimulated emission with added third level, and we left this task for future studies. It is also possible to solve this problem for initial state with two incident photons that scatters on a 3LE, but we omitted it to simplify the calculations.
IV Spontaneous decay of fully-excited transmon

### v2

@@ Abstract
We study the dynamical and spectral characteristics of a quantum three-level ladder system, interacting with a continuous electromagnetic field in one-dimensional open waveguide. Common realization of such systems is a waveguide QED setup - a superconducting artificial atom (transmon), coupled to an open microwave transmission line. We derive an analytical solution for spontaneous emission of initially excited atom, and use it to study the probability of state detection and spectral density of output photon states. We find that for strong coupling of transmon to a waveguide emitted photons show correlation in frequency and can have the same energies, even if the three-level system is anharmonic.
waveguide QED, photon-photon correlation, artificial atom, superconducting transmon, three-level system
@@ I Introduction
Waveguide quantum electrodynamics is a modern and rapidly developing field of research that studies the interaction of atoms with photons in confined geometries of waveguides (see recent reviews Roy17 and Sheremet23 ). It has a vast spectrum of applications in quantum optics and photonics, as well as in quantum technologies such as quantum communication and quantum information processing Gu17 . The adjacent field of cavity QED  — which usually focuses on manipulating quantum entanglement with atoms and photons Raimond01 , band gap engineering in condensed matter Kibis11 , nonlinear optical properties of nanostructures Savenko12 , currents induced by vacuum fluctuations Kibis13 , etc.  — together with the modern field of circuit QED Blais21 , is focused more on resonant interaction of atoms and nanostructures with single-mode fields in cavities or resonators. In contrast, waveguide QED describes the interaction of atom-like structures with a continuum of photon modes in open waveguides.
@@
Three-level systems play an important role in quantum optics and photonics, and have recently received broad attention in waveguide QED as well. Interaction of three-level emitters (3LE) with photon fields can induce some interesting effects, such as electromagnetically induced transparency (EIT) Abdumalikov10 , Autler-Townes splitting Sillanpaa09 , cross-over of these effects Anisimov11 , two-photon resonance fluorescence from experimental Gasparinetti19 and theoretical Ngaha25 point of view, giant cross-Kerr effect Hoi13a , photon-state purification Lopes24 , and many others.
Since an excited three-level system can emit or absorb two photons, it can be potentially used to induce photon-photon correlations. They can be implemented via scattering two-photon pulses on a lambda-type Roy11 or ladder-type Li15 three-level system, although a four-level system also can be used to create correlated and entangled photon pairs Zheng12 . Creation of many-photon entangled states via interaction with a $\Lambda$-type atom was studied in Ilin24 . The other way to induce photon-photon correlations is to utilize the spontaneous emission process. It can be done by driving a three-level ladder-type atom with two laser fields for each transition Chen12 , or by using a single drive on the frequency of the 0-2 transition Sathyamoorthy16 . It is also possible to create a correlated photon pair with almost instant decay of an excited 3LE for a superconducting giant-atom Gao24 , although in typical experiments for artificial atoms the process of spontaneous emission is usually of a cascaded nature Gasparinetti17 .
@@
A typical realization of a 3LE interacting with a continuum of photons in 1D waveguides is superconducting artificial atoms Kjaergaard20 , directly or side-coupled to a microwave transmission line Krantz19 . The most common type of a three-level system in this playground is a transmon qubit Koch07 . A transmon has a ladder-type structure of levels, meaning that each level can interact only with neighbouring levels. Although transmons are usually used for quantum information applications as regular qubits, i.e., two-level systems, the third level can have an important practical role in waveguide QED setups, for example to create a single-photon router Hoi11 , to characterize decoherence rates Lu21 , or to extract main parameters of qubit in open waveguide Sultanov25 . Moreover, three-level transmon structures have wide quantum optics applications, see Hoi13b and references from the previous paragraph.
@@
In this work we study the spontaneous emission of an initially excited three-level transmon, coupled to a 1D open waveguide. More traditional tasks like interaction of monochromatic wave Shen05 or photon pulse Chen11 with single qubit, or description of interaction of qubit chains with plane wave via Lippmann-Schwinger equation Fang14 or non-hermitian Hamiltonian approach Greenberg15 , or even scattering of Gaussian impulse on single qubit Greenberg23 , two-qubit system Greenberg24 , and on qubit chain Liao15 , are usually performed in the single-photon limit. In our work, however, we employ the two-excitation subspace. Our approach is based on the real-space description of photon fields. We derive an analytical solution of the non-stationary Schrödinger equation and find the probability amplitudes for the complete wavefunction of the system. Using this solution, we analyze probabilities to detect certain states of the system, and the spectral density of output two-photon fields.
For a weak radiative decay rate we find that each of the two photons has a distinct frequency, corresponding to the two transitions of the 3LE, which is in good agreement with experiments Gasparinetti17 . However, if the radiative decay is rather strong, we find that frequency correlation appears, which can result in the creation of two identical photons with equal energies. This correlation is more pronounced for low anharmonicity of the 3LE i.e., when the levels are almost equidistant. We show that the spectral density for identical photons have two maxima: on resonance with the lower two levels, or shifted by half the anharmonicity parameter. Since superconducting systems have an inherent ability to tune their internal parameters, this type of frequency correlation can be potentially used to design a source of identical photons in waveguide QED setups.
@@
The paper is organized as follows. In Section II we give a general description of the system under study and derive the main equations for the wavefunction’s amplitudes. In Section III we derive a general analytical solution in real-space. In Section IV we examine the case of a fully-excited three-level atom, both in real-space and in frequency representations. The main results are placed in this section, including analysis of state-detection probabilities, two-photon emission spectra and creation of identical photons. The conclusion of our work, as well as future prospects, are reflected in Section V. Some mathematical calculations are included in the Appendix.
@@ II Real-space description of the transmon-waveguide system
We consider a system of a ladder-type three-level emitter (3LE), located at the point $x=0$ in a one-dimensional waveguide (see figure 1(a)). In the real-space representation Shen07 and under the rotating wave approximation (RWA) it can be described by the following Hamiltonian (here and later in the text we assume that $\hbar=1$):
@@
Here states $\left|g\right\rangle,\left|e\right\rangle$ and $\left|f\right\rangle$ are the ground, first-excited, and second-excited states of the transmon. The corresponding resonant frequencies of the first and second excited states are $\Omega_{1},\Omega_{2}$, and $V_{1},V_{2}$ are their coupling constants. Energy levels and corresponding frequencies are shown in figure 1(b). $\upsilon_{g}$ is the group velocity of photons, $\delta(x)$ is the Dirac delta-function, and $a^{\dagger}(x)$ ($a(x)$) is the creation (annihilation) operator in the x-representation, which creates (destroys) a photon at position $x$ of the waveguide.
Throughout this paper we will refer to the 3LE as a transmon or an (artificial) atom, but keep in mind that our approach is universal and can be related to other physical realizations of three-level ladder-type systems.
Figure 1: (a) System under study — a three-level emitter (3LE), coupled to the center of a one-dimensional waveguide at point $x=0$. The coupling of the first and second-excited states to the waveguide are $V_{1}$ and $V_{2}$, respectively. An excited atom can emit photons through spontaneous decay, which propagate only forward due to the usage of the chiral approximation (see text for details). (b) Energy levels diagram of the bare transmon and the corresponding resonant frequencies.
The Hamiltonian (1) is written in the so-called continuous limit Blow90 , meaning that instead of a discrete summation over $x$ (or over modes $k$), we have a continuous integration. In this limit, the commutation relation between photon operators is as follows:
@@
hence the photon operators, usually dimensionless for discrete modes, in the continuous limit have the dimension of the inverse square root of coordinate, $a(x)\sim 1/\sqrt{x}$. The dimension of the coupling constants $V_{1},V_{2}$, correspondingly, is $V\sim\omega\sqrt{x}$.
The first two terms in (1) correspond to a free atom and a free field, and the following two terms describe the interaction of photon fields with the first and second excited states of the transmon, respectively. Note that there is no direct transition between the ground and the second-excited state $\left|f\right\rangle\leftrightarrow\left|g\right\rangle$. Hamiltonian (1) is practically the Jaynes-Cummings Hamiltonian, extended for a three-level system, and it conserves the number of excitations in the system. In this work, we adhere to two excitations due to the two possible transitions of the transmon. We should note that in real superconducting waveguides, there exist two dissipation sources which are usually described by the non-radiative decay rate $\Gamma_{nr}$ and the decay rate to non-guided modes $\Gamma_{ng}$. However, they are much less than $\Gamma_{1D}$, the decay rate to the waveguide guided mode. Therefore, the coupling efficiency $\beta=\Gamma_{1D}/(\Gamma_{1D}+\Gamma_{nr}+\Gamma_{ng})$ is close to unity. For superconducting systems it can reach values of $\beta\approx 0.999$ Sheremet23 , so our exclusion of dissipation is well justified.
Since our atom has three levels, the wave function should be written in the two-excitation subspace as follows:
@@
where the first probability amplitude $\alpha(t)$ corresponds to a excited transmon in state $\left|f\right\rangle$ and no photons in the waveguide, the second amplitude $\beta(x,t)$ corresponds to a transmon excited to the $\left|e\right\rangle$ state and one photon at position $x$, and the final amplitude $\gamma(x_{1},x_{2},t)$ contains a non-excited atom and two photons at positions $x_{1}$ and $x_{2}$. State $\left|0\right\rangle$ designates the vacuum Fock state. The coefficient $1/\sqrt{2}$ in the last term is a normalization constant for a two-photon state in the Fock basis.
Because of the continuous limit Blow90 , probability amplitudes with photons have the dimension of the inverse square root of coordinate: $\beta(x,t)\sim 1/\sqrt{x}$, and $\gamma(x_{1},x_{2},t)\sim 1/x$ due to the presence of two photons.
We should also note that the wavefunction (3), as well as the Hamiltonian (1), describe photon propagation only in the forward direction. This is the so-called chiral approximation, and it can be performed by separating the Hamiltonian for odd and even modes Shen07 . Chiral, or non-reciprocal, waveguides are realized by creating a suppression of photon propagation in the backward direction and can be implemented in current waveguide QED experimental setups Guimond20 . However, we are mainly interested in the case of spontaneous decay of an initially excited 3LE, which should be symmetrical with respect to the 3LE placement, and thus bidirectionality has little effect on our system. So, in order to simplify the analytical derivation, we choose to consider only forward propagation.
Using the standard non-stationary Schrödinger equation $i\partial_{t}\left|\Psi\right\rangle=H\left|\Psi\right\rangle$ with the Hamiltonian (1) and the wave function (3), we derive the equations for all probability amplitudes:
@@
In the last two equations, we use the fact that photons are fundamentally indistinguishable, which in our case can be stated as $\gamma(x_{1},x_{2},t)=\gamma(x_{2},x_{1},t)$. All further calculations are based on equations (4-6).
@@ III Probability amplitudes in real-space
To find explicit expressions for all three amplitudes of the wave function, we use a method incorporated in Mukhopadhyay24 for the calculation of quantum Rabi oscillation dynamics and also employed in Lopes24 for photon state purification via 3LE interaction. We have already used this approach in Chuikin24 for the description of stimulated emission in waveguide QED.
First, let’s define a Fourier transform for amplitudes with a photonic part, which allows us to switch from real-space to k-space of frequency modes:
@@
Inverse transformation is done by changing the sign of the exponents.
Here and further we assume that the dispersion is linear, $k=\omega/\upsilon_{g}$, which is well justified if the atom and photon resonant frequencies are far from the cut-off frequency of the waveguide Shen09 .
By using (7) in (6) together with the integral representation of the delta function as $\delta(x)=1/2\pi\int dk\;e^{ikx}$, we obtain an equation for the two-photon amplitude in k-space:
$\begin{gathered}\frac{d}{dt}\gamma(k_{1},k_{2},t)+i\upsilon_{g}(k_{1}+k_{2})\gamma(k_{1},k_{2},t)=\\ =-\frac{i}{\sqrt{2}\sqrt{2\pi}}V_{1}{e^{-i\Omega_{1}t}}\left[{\beta(k_{2},t)+\beta(k_{1},t)}\right].\end{gathered}$ (8)
@@
where we assume that at the initial time $t=0$ no two-photon states are present. Transforming this expression back to real-space using the inverse of (7), we get:
@@
where we introduced a Heaviside step-function, $\theta(x>0)=1,\;\theta(x<0)=0$. In order to avoid a jump at $x=0$, we apply the standard regulation $\theta(0)=1/2$.
The next thing we need is to find $\beta(x,t)$, which then can be used to determine the two-photon amplitude (10). Taking $\gamma(x_{1},x_{2},t)$ for $x_{2}=0$ and substituting it into (5), we obtain:
@@
where we define the decay rate of the first-excited level as $\Gamma_{1}=V_{1}^{2}/\upsilon_{g}$, as well as some additional parameter:
@@
To find an explicit expression for $\beta(x,t)$, we use the same procedure as for $\gamma(x_{1},x_{2},t)$. Multiplying (11) by $e^{-ikx}/\sqrt{2\pi}$ and integrating over $x$, we can use the Fourier transform (7) to rewrite equation (11) in k-space. Then the formal solution can be found as:
@@
where $\beta_{0}(k)=\beta(k,0)$ is an amplitude for an incident photon with frequency $\omega=k\upsilon_{g}$ at the initial time $t=0$ and $F(k,t)$ is the Fourier transform of coefficient (12). Translating (13) back to real-space, we get:
@@
where we introduce the frequency between levels $\left|e\right\rangle$ and $\left|f\right\rangle$ as $\Delta_{\Omega}=\Omega_{2}-\Omega_{1}$ (see figure 1(b)). The change of the upper limit in the last integral is due to the step function $\theta\left(x-\upsilon_{g}t^{\prime}\right)$.
Expression (14), aside from the amplitude $\alpha(t-x/\upsilon_{g})$, also contains the shifted value $\beta(\upsilon_{g}t^{\prime}-x,t-x/\upsilon_{g})$, so it can’t be considered explicit. However, by performing a corresponding shift in (14), one can show that the recursive part will be cut off due to the Heaviside functions, and the remaining part will contain only terms proportional to the incident single-photon field $\beta_{0}$ and the amplitude of the second-excited state $\alpha(t)$. Detailed calculations can be found in Appendix A. Thus, we obtain the explicit expression for the single-photon amplitude:
@@
By taking (15) at point $x=0$, we can easily find the remaining amplitude of the second-excited state from (4):
$\alpha(t)=\alpha(0)e^{-\frac{\Gamma_{2}}{2}t}-iV_{2}e^{-\frac{\Gamma_{2}}{2}t}\int\limits_{0}^{t}d\tau e^{\left({i\Delta_{\Omega}-\frac{\Gamma_{1}}{2}+\frac{\Gamma_{2}}{2}}\right)\tau}\beta_{0}(-\upsilon_{g}\tau),$ (16)
where $\alpha(0)$ corresponds to the initial state with a fully excited transmon, and $\Gamma_{2}=V_{2}^{2}/\upsilon_{g}$ is the decay rate of the second-excited state, similar to $\Gamma_{1}$.
The obtained expressions (10), (15) and (16) are the key results of this section. Given a particular initial state, we can derive an explicit analytic expression for the full wave function (3). These equations admit two possible initial conditions: 1) The transmon is fully excited in state $\left|f\right\rangle$ and there are no free photons in the waveguide; 2) The transmon is excited to state $\left|e\right\rangle$ and there is a single incident photon with an arbitrary line shape $\beta_{0}\equiv\beta(k,0)$. The first condition describes spontaneous emission of a fully excited three-level system, and we analyze it in detail below. The second one describes the process of stimulated emission with an added third level, and we leave this task for future studies. It is also possible to solve this problem for an initial state with two incident photons scattering on a 3LE, but we omitted it to simplify the calculations.
IV Spontaneous decay of the fully-excited transmon